If Japan must give up the production of 75 computers to produce 25 additional cellular telephones, the opportunity cost of producing 4 cellular telephones is _____ computers.

Select one:

a. 12

b. 79

c. 71

d. 4
Answer #1

Answer

The correct answer is (a) 12

Assuming that There is rate of substitution is constant, If Japan must give up the production of 75 computers to produce 25 additional cellular telephones, then In order to 1 cellular telephone he has to give up (75/25) = 3 computers.

This implies In order to produce 4 cellular phone he has to sacrifice 3*4 = 12 computers.

Opportunity cost is the value you next best alternative that you have to sacrifice in order to take the alternative you have chosen.

Here, In order to produce 4 cellular phones he has to sacrifice 3*4 = 12 computers

So, Opportunity cost of producing 4 cellular phones in 12 computers.

Hence, the correct answer is (a) 12